### v2

# Melting curve of correlated iron at Earth’s core conditions from machine‑learned DFT+DMFT

Abstract
Reliable constraints on iron’s melting curve at Earth’s inner‑core boundary require accurate finite‑temperature electronic correlations, yet DFT+DMFT calculations remain too costly for large‑scale thermodynamic sampling. Here, we develop a machine‑learning accelerator for charge self-consistent DFT+DMFT by training E(3)-equivariant graph neural networks to predict the local self-energy and Fermi level from atomic environments, providing an efficient warm start to the DMFT self-consistency loop. Using high‑throughput data for Fe, FeO, and NiO, we obtain a 2–4 times reuduction in DMFT iterations. Leveraging this improvement, we generate correlated energies and forces for Fe at core pressures, train a neural‑network interatomic potential, and determine the melting curve via two‑phase coexistence simulations. We obtain a predicted melting temperature of 6225 K at 330 GPa.
Correlated electron systems occupy a central place in modern condensed-matter physics. Their rich behaviors, including metal–insulator transitions [mitAbrahams1996, mitMasatoshi1998], unconventional magnetism [ucmBroholm2020, ucmLee2013], and high-temperature superconductivity [htsDagotto1994, htsKeimer2015, htsStewart2017], arise from strong local interactions, particularly in transition-metal compounds where lattice, spin, orbital, and charge degrees of freedom are tightly intertwined [tmoAnisimov1997, tmoTokura2000]. Understanding how this interplay governs electronic properties across pressures, temperatures, and compositions remains one of the major challenges in computational materials physics and is essential for problems spanning functional oxides to materials under extreme pressure and temperature conditions.
Fe at the conditions of Earth’s core exemplifies the challenges of simulating correlated systems [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015]. The melting curve of Fe at hundreds of gigapascals controls the temperature at the inner-core boundary (ICB) and thus constrains the thermal history and dynamics of our planet, yet experimental determinations at these pressures are extremely challenging and theoretical predictions span a broad range of temperatures. While recent experiments have begun to converge on values near 6200 K at ICB pressure [emcAnzellini2013, emcKraus2022, emcBalugani2024], theoretical predictions from density functional theory (DFT) calculations remain scattered, with discrepancies often exceeding 1000 K [aimcAlfe2002, aimcAlfe2009, aimcBelonoshko2021, aimcBelonosko2000, aimcGonzalez2023, aimcSola2009, aimcStixrude2014, aimcSun2022, aimcSun2023, aimcWu2024]. A fundamental limitation of standard DFT approaches is their inadequate treatment of dynamical correlations in Fe [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015]. Under core conditions, hexagonal close-packed (hcp) Fe and body centered cubic (bcc) Fe retain their Fermi-liquid character while fcc Fe displays markedly non-Fermi liquid behavior, with relative free energy differences changing significantly with temperature [ieccPourovskii2019]. Incorporation of these correlation-induced differences are neglected with conventional exchange-correlation functionals, but must be included for accurate simulations of liquid phases.
Dynamical mean-field theory (DMFT), which captures local correlations in a fully dynamical and non-perturbative manner [dmftKotliar2006], offers a path forward. Coupled to DFT, it allows for realistic simulations of correlated materials by taking into account effects such as quasiparticle renormalizations, and provides access to electronic and thermodynamic properties such as spectral functions and finite-temperature phase diagrams [dmftArpita2019]. However, this power comes at a substantial computational cost. Each DMFT calculation requires repeated solution of the quantum impurity problem until impurity and lattice occupations become consistent. Determining a melting curve requires sampling thousands of large-scale atomic configurations through molecular dynamics, a task that has remained prohibitively expensive within DFT+DMFT.
Recently, machine learning has begun to be explored as a way to reduce the cost of DMFT calculations, for example through surrogate impurity solvers, learned corrections to approximate Green’s functions, and data-driven models trained on simplified parameter spaces [mlgfMitra2025, mlgfDong2024, mlgfSturm2021, mlgfArsenault2014]. While these studies demonstrate that key DMFT quantities can often be approximated efficiently in selected settings, most existing approaches have been confined to model Hamiltonians or narrow material classes, and the learned targets are not always constructed to respect the analytic structure and symmetry constraints required in realistic DFT+DMFT workflows. It therefore remains unclear whether machine learning can provide a robust, physics-consistent acceleration of DFT+DMFT for correlated solids with the accuracy needed for thermodynamic predictions, particularly under the extreme conditions relevant to Earth’s core.
In this work, we demonstrate that embedding physical constraints into machine-learned self-energies enables, for the first time, systematic DFT+DMFT determination of the iron melting curve at core conditions. Our approach exploits the analytic structure of the self-energy: we decompose it into a static high-frequency limit $\Sigma(\infty)$ and a frequency dependent term $\Sigma(i\omega)$, represented compactly through Legendre coefficients, with each component predicted by E(3)-equivariant graph neural networks that respect the point-group symmetry of the local atomic environment. This physics-informed representation reduces the learning problem to smooth, real-valued targets while preserving the causal structure essential for DMFT self-consistency [mlgfArsenault2014].
The resulting model accelerates DMFT convergence by a factor of two to four across Fe, FeO, and NiO, enabling high-throughput generation of correlated electronic structure data. We leverage this capability to train a machine learned interatomic potential (MLIP) for iron and perform two-phase coexistence simulations spanning 307–352 GPa. Our predicted melting temperature of 6225K at 330 GPa agrees well with the recent experimental measurements [emcAnzellini2013, emcKraus2022, emcBalugani2024], providing strong validation that large-scale molecular dynamics (MD) simulations using DMFT-based MLIPs are viable for phase-equilibrium calculations in planetary interiors.
Within DFT+DMFT, electronic correlations enter entirely through the local self-energy. For each correlated atom and orbital, the DMFT loop determines a frequency-dependent self-energy $\Sigma(i\omega_{n})$ that, together with the noninteracting lattice Green’s function, gives the electronic structure and thermodynamic properties of the crystal. The real part of $\Sigma$ controls band renormalization and shifts, while the imaginary part governs lifetimes, scattering rates, and the appearance of incoherent spectral weight. In practice, for every crystal structure and thermodynamic state point $(T,P)$ one must find a fixed point of the DMFT equations at which the impurity and lattice quantities (such as occupancies and local Green’s functions) become consistent, yielding a converged self-energy $\Sigma$.
Formally, this converged self-energy is a functional of the local environment: it depends on the atomic configuration, the chemical composition, and the thermodynamic parameters, as well as on the choice of correlated subspace and interaction parameters. In this work we make this viewpoint explicit and seek an approximate, parameterized functional $\Sigma^{\mathrm{ML}}[{\mathcal{S}}](i\omega_{n})$ that maps a given structure $\mathcal{S}$} (including atomic positions, species, and volume) to an orbital-resolved local self-energy close to the DMFT fixed point. The goal is not to replace DMFT entirely, but to provide an initial approximation $\Sigma^{\mathrm{ML}}$ that lies close enough to $\Sigma$ such that this fixed point may be reached more rapidly.
A naive strategy would be to train a neural network to output $\Sigma(i\omega_{n})$ directly on a discrete Matsubara frequency grid. This is undesirable for several reasons. First, the DMFT self-energy is complex-valued and, when obtained from quantum Monte Carlo impurity solvers, can be noisy at high frequencies. Second, $\Sigma(i\omega_{n})$ obeys strict analytic constraints: it is a causal function with a well-defined high-frequency expansion, $\Sigma(i\omega_{n})=\Sigma(\infty)+{\cal O}(1/i\omega_{n})$ [sigmaWang2011], and its imaginary part has definite sign on the imaginary axis. Direct regression on a large set of complex values risks violating these constraints unless they are enforced explicitly.
Instead, we adopt a physics-constrained representation tailored to the analytic structure of the self-energy. The real part of the self-energy approaches a frequency-independent limit as $\omega_{n}\rightarrow\infty$:
$Re\left\{\Sigma(i\omega_{n})\right\}\xrightarrow{\omega_{n}\to\infty}\Sigma(\infty),$ (1)
while the imaginary part decays to zero. $\Sigma(\infty)$ captures the static Hartree-Fock contribution from local interactions. We therefore decompose the self-energy as
$\Sigma(i\omega_{n})=\Sigma(\infty)+\Delta\Sigma(i\omega_{n}),$ (2)
where $\Delta\Sigma(i\omega_{n})\to 0$ at high frequency. This separation isolates the dynamical correlations in $\Delta\Sigma$ from the static component $\Sigma(\infty)$, each of which can be learned independently.
The dynamical part $\Delta\Sigma(i\omega)$ is then Fourier transformed to imaginary time and expanded in a Legendre polynomial basis [Boehnke2011],
$\Delta\Sigma(\tau)=\sum_{\ell=0}^{\ell_{\max}}\sqrt{2\ell+1}\,P_{\ell}(x(\tau))\,\Sigma_{\ell},$ (3)
where $x(\tau)=2\tau/\beta-1$ and $P_{\ell}$ are Legendre polynomials. The coefficients $\Sigma_{\ell}$ decay rapidly with $\ell$, providing a compact representation. For example, we find $\ell_{\mathrm{max}}=30$ sufficient for Fe and FeO at 5000 K, while NiO at 611 K requires $\ell_{\mathrm{max}}=70$ due to finer structure at lower temperature.
In this representation, the Legendre expansion is designed to enforce that $\Delta\Sigma(i\omega)$ is a smooth function on the Matsubara frequency axis and yields a self-energy with the proper high-frequency properties. Because the Legendre coefficients are real, the learning problem is reduced to predicting a small set of real numbers per orbital and site rather than a noisy complex function on a dense grid. Therefore, we are learning a compact, physically motivated set of coefficients that encode the self-energy under known analytic constraints.
In addition to the self-energy, DMFT calculations require knowledge of the Fermi level $E_{f}$ to enforce the total-charge constraint. Small errors in $E_{f}$ can lead to discrepancies between impurity and lattice occupancies and thus slow or even prevent convergence. We therefore include $E_{f}$ as an additional output of our physics-informed functional.
Given the predicted $\Sigma^{\mathrm{ML}}(\infty)$ and Legendre coefficients, the approximate self-energy on the Matsubara axis is reconstructed as
$\Sigma^{\mathrm{ML}}(i\omega_{n})=\Sigma^{\mathrm{ML}}(\infty)+\Delta\Sigma^{\mathrm{ML}}(i\omega_{n}),$ (4)
where $\Delta\Sigma^{\mathrm{ML}}(i\omega_{n})$ is obtained by transforming the Legendre expansion back to the Matsubara axis. Both forward and reverse transformations are performed by the TRIQS software package [Parcollet2015, Boehnke2011]. This $\Sigma^{\mathrm{ML}}$, together with the predicted $E_{f}$, provides the initial condition for the DMFT self-consistency loop.
Including $E_{f}$ as a learned quantity is an essential for achieving substantial speedups. In DFT+DMFT the Fermi level is normally adjusted iteratively so that the total electron number satisfies the charge constraint, and poor initial guesses can severely delay convergence. By treating $E_{f}$ on the same footing as $\Sigma(\infty)$ and the Legendre coefficients, we define a joint functional $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ that maps a crystal structure to a complete, physics-consistent set of input parameters.
To implement this functional, we represent each crystal structure as a periodic atomic graph and approximate $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ with an $E(3)$-equivariant graph neural network built using the open-source e3nn software package [e3nnGieger2022, e3nnKondor2018, e3nnThomas2018, e3nnWeiler2018]. Atoms are nodes, connected to neighbors within a radial cutoff, and are encoded by learned embeddings of their chemical species and mass. Relative position vectors between neighbors are expanded in spherical harmonics and radial basis functions, so that all edge features and intermediate representations transform as irreducible representations of $E(3)$. A small stack of equivariant message-passing layers then builds a symmetry-respecting description of each local environment that is translationally invariant and equivariant under rotations and inversion. Further architectural details are given in the Supplemental Material.
From the final equivariant representation, we construct scalar invariants for prediction. For each inequivalent correlated atom, orbital-specific readout heads map the local features to the static component $\Sigma(\infty)$ and the truncated set of Legendre coefficients $\Sigma_{\ell}$ for that orbital. A separate global readout, obtained by pooling scalar channels over all atoms, yields the Fermi level $E_{f}^{\mathrm{ML}}$. Therefore, a single forward pass of the network maps a crystal structure $\mathcal{S}$ to $\Sigma^{\mathrm{ML}}(\infty)$, $\{\Sigma_{\ell}^{\mathrm{ML}}\}$, and $E_{f}^{\mathrm{ML}}$; these are then used to reconstruct $\Sigma^{\mathrm{ML}}(i\omega_{n})$ and provide a fully physics-consistent warm start for the subsequent DFT+DMFT self-consistency cycle.
To assess the accuracy of the physics-informed functional, we first compare its predictions for the self-energy against fully converged DFT+DMFT results on held-out test structures. For Fe and FeO, structural configurations were generated from MD simulations of high‑symmetry crystalline prototypes (e.g., hcp/bcc/fcc for Fe, B2 phase for FeO) to ensure coverage of both ordered and disordered environments. 2.5‑ps DFT‑MD trajectories (0.5 fs timestep) of 4–8 atom cells were generated and representative snapshots were selected by k‑medoids clustering in structural fingerprint space [fpZhu2016], yielding 500 configurations in total (100-200 per trajectory). Since the fingerprint represents the chemical environment of each crystal structure, the fingerprint distance allows us to quantify the overall difference in chemical environments between any two crystal structures. This ensures that we obtain diverse chemical environments for the solver to “learn” in an attempt to aid generalization of results to unseen structures. For NiO we use configurations from our previous study [Rao2024]. Full details of dataset generation and selection are provided in the Supplemental Material. We reserve 10% as a held-out test set.
In Fig. 1(a-c), we show the self-energy predictions for 4 random test structures as compared to the self-energies from DFT+DMFT. We find good agreement between predicted and actual self-energies. The errors in the $\Sigma(i\omega)$ predictions for Fe and FeO were considerably lower than that for NiO. This is likely due to the much lower temperature calculations for NiO, which subsequently lowers Matsubara axis spacing and increases the density of points on the grid.
We also present the errors from the $\Sigma(\infty)$ predictions as well as the predictions of the Fermi level $E_{f}$ in table 1. We find in this case, it is instead NiO that provides the lowest error. This could be due to the greater number of training points available for NiO, as well as the similarity in the structures generated as these were sampled directly from perturbations of the atomic structures of certain symmetries. The predictions are reasonable for each compound however, especially as a starting guess for the DFT+DMFT loop. Since further data generation only improves the predictions, one can set up an active learning loop to iteratively improve prediction accuracy.
MAE (eV) $E_{f}$ $\Sigma_{z^{2}}$ $\Sigma_{x^{2}-y^{2}}$ $\Sigma_{xz}$ $\Sigma_{yz}$ $\Sigma_{xy}$
Fe 0.217 0.302 0.298 0.315 0.310 0.319
FeO 0.707 0.458 0.341 0.417 0.385 0.408
NiO 0.208 0.198 0.200 0.378 0.3873 0.314
Table 1: Mean absolute errors of predictions for $E_{f}$ and $\Sigma(\infty)$ orbital components on test set
Figure 1: Top row: Predicted real (red) and imaginary (blue) parts of the self energy vs DMFT (black) self-energies for 4 randomly selected structures in test sets of Fe, FeO and NiO. Bottom row: Convergence for Fe, FeO and NiO versus number of DFT+DMFT iterations.
Figure 1 (d-f) shows the convergence of impurity-lattice occupation difference $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|$ as a function of DMFT iteration for Fe, FeO, and NiO, averaged over the data for 20 randomly chosen structures. With zero initialization, Fe requires approximately 9 iterations to reach our convergence criterion of $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|<0.05$; with ML-predicted self-energies, convergence is achieved in 2–3 iterations. Similar speedups are observed for FeO (9 vs. 5 iterations) and NiO (6 vs. 2 iterations). Since each CTQMC iteration dominates the computational cost, this translates directly into a factor of 2–4 reduction in wall-clock time per structure. The learned functional thus acts as a nearly cost-free preconditioner for the DMFT loop, turning the self-energy from a repeated, expensive output into a one-shot, physics-informed initial condition.
Leveraging this acceleration, we performed high-throughput DFT+DMFT calculations on an additional 600 Fe configurations at pressures and temperatures relevant to Earth’s core, increasing the Fe dataset to 1100 configurations with correlated energies and forces. Using all Fe data collected so far, we trained a NequIP-based machine-learned interatomic potential (MLIP) [nequipTan2025, nequipBatzner2022], achieving test-set mean absolute errors of 69.2 meV/atom for the energy and 76.7 meV/Å for the forces. Because the reference energies and forces are obtained from a CTQMC impurity solver, the training labels carry finite statistical uncertainty [huangMachineLearningDiffusion2023]; accordingly, the reported test errors reflect both model discrepancy and an intrinsic noise floor. In the high-temperature regime of interest (e.g., $T=5000$ K, $k_{B}T\approx 430$ meV), the resulting per-atom energy error remains sub-$k_{B}T$, enabling stable large-scale MD and solid–liquid coexistence simulations. To assess sensitivity to the electronic temperature used for DFT+DMFT labeling, we recomputed the electronic free energies for representative Fe configurations at T = 5000 and 6000 K and find changes in relative free-energy differences of $\leq$ 6 meV/atom, well below the MLIP test errors.
We determined melting temperatures using the solid–liquid coexistence method [aimcAlfe2009], which avoids the hysteresis inherent in single-phase superheating or supercooling approaches. An elongated supercell (9216 atoms) containing coexisting solid (hcp) and liquid regions is equilibrated within the NVE ensemble; if the temperature lies above $T_{m}$, the solid region melts, whereas below $T_{m}$ the liquid crystallizes. The phase fractions were monitored using local bond-order analysis [pfaLechner2008]. Figure 2 illustrates a representative coexistence configuration at 307.8 GPa: the density profile along the long axis exhibits pronounced oscillations in the crystalline region, characteristic of atomic layering, and a comparatively smooth profile in the liquid region, confirming a well-defined solid–liquid interface. The temperature trace in the inset shows only moderate fluctuations after 15ps during the NVE run, indicating stable coexistence conditions. Simulations were thermalized for 8 ps and subsequently evolved for 20 ps or 30 ps, depending on phase stabilization, with a 0.5 fs timestep. Pressure was sampled from finite-difference stress calculations including the kinetic contribution, and averaged over the final 5 ps.
Figure 2: Density profile along the long axis of simulation cell taken at 20 ps into equilibration for simulation at 307.8 GPa. Shown in the insets are a snapshot of the atomic configuration at 20 ps as well as the temperature variation over the length of the simulation, which stabilizes around 15 ps. The 2 coexistence interfaces can been seen on both sides of the liquid.
Our melting results, spanning 307–352 GPa, is shown in Fig. 3 alongside experimental determinations and prior ab initio predictions. The results are well described by a Simon-Glatzel relation [simonglatz],
$T_{m}=T_{0}\left(\frac{P-P_{0}}{a}+1\right)^{1/c},$ (5)
with $T_{0}=6167$ K, $P_{0}=323.1$ GPa, $a=298.73$ GPa, and $c=2.43$. At the inner-core boundary pressure of 330 GPa, we obtain $T_{m}=6225\pm 42$ K, in excellent agreement with the laser-heated diamond anvil cell measurements of Anzellini et al. ($6230\pm 500$ K) [emcAnzellini2013] and the ramp-compression experiments of Kraus et al. [emcKraus2022]. It is also compatible with the recent shock-compression XAS constraints of Balugani et al., who extrapolate an upper bound of 6202 ± 514 K for Fe at 330 GPa [emcBalugani2024]. Our result lies above the resistance-heated DAC determination of Sinmyo et al. ($5500\pm 220$ K) [emcSinmyo2019]; however, as noted in Ref. [emcAnzellini2013], different melting criteria in static experiments can yield systematically different temperatures, and the discrepancy among experiments themselves remains an active area of investigation.
Our agreement with recent experimental constraints also provides insight into why melting predictions for Fe at core pressures have remained so dispersed. Most published ab initio melting curves based on standard DFT span a broad range and are frequently higher than experimental estimates, with discrepancies reaching several hundred kelvin and an overall scatter approaching $10^{3}$ K [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009]. This sensitivity likely reflects, in part, that commonly used high-$T$ DFT melting workflows dynamical correlations only approximately, thereby missing high-temperature spin fluctuations that contribute to the free energy [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015].
Fe at these conditions is paramagnetic, yet thermal spin fluctuations can sustain sizable instantaneous local moments and associated entropy contributions [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015]. If these magnetic/electronic excitations are more pronounced in the liquid than in the solid, they increase the entropy of melting and reduce the coexistence temperature through the equilibrium condition $\Delta G(P,T_{m})=0$, i.e. $T_{m}=\Delta H/\Delta S$.
DFT+DMFT addresses this limitation by treating the $3d$ correlations dynamically through the frequency-dependent self-energy $\Sigma(i\omega_{n})$, capturing quasiparticle renormalization and fluctuating local moments on equal footing. That our melting curve falls within experimental uncertainty while using interaction parameters $(U,J)$ adopted from prior DFT+DMFT studies [ieccPourovskii2017, chsGavriliuk2023, chsHo2024] (rather than fitted to melting data) indicates genuine predictive power and is consistent with the view that dynamical correlations account for a substantial part of the long-standing discrepancy between simulation and experiment for Fe at core conditions.
Figure 3: Full melting curve of Iron as predicted by our 2-phase simulations and compared to data from various experimental [emcAnzellini2013, emcSinmyo2019, emcLi2020, emcKraus2022, emcBalugani2024] computational studies [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009].
In conclusion, we introduced a physics‑informed machine‑learning approach that substantially reduces the computational cost of realistic DFT+DMFT calculations. Rather than directly fitting the noisy Matsubara self‑energy as an unconstrained output, we learn a compact representation built from the high‑frequency limit and a truncated set of real Legendre coefficients, together with the Fermi level. Implemented with E(3)‑equivariant graph neural networks, this provides a physically consistent warm start to the DMFT fixed point and reduces the number of costly impurity iterations by a factor of two to four across Fe, FeO, and NiO. This acceleration turns DFT+DMFT into a practical engine for generating correlated energies and forces at scale. Using the accelerated workflow, we expanded our Fe dataset and trained a NequIP‑based MLIP that enables large‑scale molecular dynamics for phase equilibrium. Applying solid–liquid coexistence simulations in the 307–352 GPa range, we obtained the melting curve of Fe under core conditions and predict $T_{m}=6225$ K at 330 GPa, in close agreement with recent experimental constraints. More broadly, this approach turns DFT+DMFT into a scalable engine for finite‑temperature simulations of correlated materials, with applications ranging from transition‑metal compounds to matter under planetary interior conditions.
Acknowledgements—This work was supported by the startup funds of the office of the Dean of SASN of Rutgers University-Newark. The authors acknowledge the Office of Advanced Research Computing (OARC) at Rutgers for providing access to the Amarel cluster and associated research computing resources.
Data Availability—The software used in prediction of the self-energies and Fermi level has been made available at https://github.com/Rutgers-ZRG/SigML

### v3

@@
-Melting curve of correlated iron at Earth’s core conditions from machine‑learned DFT+DMFT
+Physics-Constrained Self-Energy Warm Starts for Charge-Self-Consistent DFT+DMFT: Application to Iron at Core Conditions
 Abstract
-Reliable constraints on iron’s melting curve at Earth’s inner‑core boundary require accurate finite‑temperature electronic correlations, yet DFT+DMFT calculations remain too costly for large‑scale thermodynamic sampling.
-Here, we develop a machine‑learning accelerator for charge self-consistent DFT+DMFT by training E(3)-equivariant graph neural networks to predict the local self-energy and Fermi level from atomic environments, providing an efficient warm start to the DMFT self-consistency loop.
-Using high‑throughput data for Fe, FeO, and NiO, we obtain a 2–4 times reuduction in DMFT iterations.
-Leveraging this improvement, we generate correlated energies and forces for Fe at core pressures, train a neural‑network interatomic potential, and determine the melting curve via two‑phase coexistence simulations.
-We obtain a predicted melting temperature of 6225 K at 330 GPa.
+Charge self-consistent DFT+DMFT quantitatively captures dynamical electronic correlations in real materials, but its cost precludes the large-scale thermodynamic sampling required for phase boundaries and equations of state.
+Here, we develop a physics-constrained machine-learning warm start for realistic DFT+DMFT: E(3)-equivariant graph neural networks predict a compact, real-valued representation of the local self-energy and Fermi level—{ $\Sigma(\infty),\,\Sigma_{\ell},\,E_{f}\,$}—tied to the known high-frequency and analytic structure of $\Sigma(i\omega_{n})$, and used to initialize the full DFT+DMFT self-consistency cycle.
+Across metallic Fe, correlated FeO, and Mott-insulating NiO, the scheme yields a 2–4 times reduction in the number of DMFT iterations required to reach self-consistency.
+As a demanding application, we leverage this capability to generate correlated energies and forces for Fe at core pressures, train an equivariant machine-learned interatomic potential, and determine the hcp-Fe melting curve by solid–liquid coexistence simulations in the NVE ensemble in 9216-atom cells.
+We obtain a melting temperature of 6225 K at 330 GPa, in agreement with recent experimental constraints and consistent with the view that dynamical electronic correlations contribute to the discrepancy between DFT-based predictions and experiment.
 Correlated electron systems occupy a central place in modern condensed-matter physics.
-Their rich behaviors, including metal–insulator transitions [mitAbrahams1996, mitMasatoshi1998], unconventional magnetism [ucmBroholm2020, ucmLee2013], and high-temperature superconductivity [htsDagotto1994, htsKeimer2015, htsStewart2017], arise from strong local interactions, particularly in transition-metal compounds where lattice, spin, orbital, and charge degrees of freedom are tightly intertwined [tmoAnisimov1997, tmoTokura2000].
+Their rich behaviors, including metal–insulator transitions [1, 35], unconventional magnetism [14, 47], and high-temperature superconductivity [16, 36, 64], arise from strong local interactions, particularly in transition-metal compounds where lattice, spin, orbital, and charge degrees of freedom are tightly intertwined [4, 72].
 Understanding how this interplay governs electronic properties across pressures, temperatures, and compositions remains one of the major challenges in computational materials physics and is essential for problems spanning functional oxides to materials under extreme pressure and temperature conditions.
-Fe at the conditions of Earth’s core exemplifies the challenges of simulating correlated systems [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015].
+Fe at the conditions of Earth’s core exemplifies the challenges of simulating correlated systems [57, 58, 74].
 The melting curve of Fe at hundreds of gigapascals controls the temperature at the inner-core boundary (ICB) and thus constrains the thermal history and dynamics of our planet, yet experimental determinations at these pressures are extremely challenging and theoretical predictions span a broad range of temperatures.
-While recent experiments have begun to converge on values near 6200 K at ICB pressure [emcAnzellini2013, emcKraus2022, emcBalugani2024], theoretical predictions from density functional theory (DFT) calculations remain scattered, with discrepancies often exceeding 1000 K [aimcAlfe2002, aimcAlfe2009, aimcBelonoshko2021, aimcBelonosko2000, aimcGonzalez2023, aimcSola2009, aimcStixrude2014, aimcSun2022, aimcSun2023, aimcWu2024].
-A fundamental limitation of standard DFT approaches is their inadequate treatment of dynamical correlations in Fe [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015].
-Under core conditions, hexagonal close-packed (hcp) Fe and body centered cubic (bcc) Fe retain their Fermi-liquid character while fcc Fe displays markedly non-Fermi liquid behavior, with relative free energy differences changing significantly with temperature [ieccPourovskii2019].
-Incorporation of these correlation-induced differences are neglected with conventional exchange-correlation functionals, but must be included for accurate simulations of liquid phases.
-Dynamical mean-field theory (DMFT), which captures local correlations in a fully dynamical and non-perturbative manner [dmftKotliar2006], offers a path forward.
-Coupled to DFT, it allows for realistic simulations of correlated materials by taking into account effects such as quasiparticle renormalizations, and provides access to electronic and thermodynamic properties such as spectral functions and finite-temperature phase diagrams [dmftArpita2019].
+While recent experiments have begun to converge on values near 6200 K at ICB pressure [5, 40, 7], direct measurement at the ICB itself remains inaccessible: laser-heated diamond-anvil cell data extrapolate upward from $\sim$200 GPa, while ramp-compression data extrapolate downward from $\sim$600 GPa, so the quoted experimental uncertainties already reflect extrapolation in addition to measurement error.
+The phase of Fe at melting under these conditions is also an open experimental question, with a recent preprint reporting bcc formation near melting above 200 GPa [38].
+Theoretical predictions from density functional theory (DFT) calculations remain scattered, with discrepancies often exceeding 1000 K [2, 3, 10, 9, 26, 63, 65, 68, 67, 79].
+A fundamental limitation of standard DFT approaches is their inadequate treatment of dynamical correlations in Fe [57, 58, 74].
+Under core conditions, hexagonal close-packed (hcp) Fe and body centered cubic (bcc) Fe retain their Fermi-liquid character while fcc Fe displays markedly non-Fermi liquid behavior, with relative free energy differences changing significantly with temperature [58].
+These correlation-induced differences are neglected by conventional exchange-correlation functionals but must be included for accurate simulations of liquid phases.
+Dynamical mean-field theory (DMFT), which captures local correlations in a fully dynamical and non-perturbative manner [39], offers a path forward.
+Coupled to DFT, it allows for realistic simulations of correlated materials by taking into account effects such as quasiparticle renormalizations, and provides access to electronic and thermodynamic properties such as spectral functions and finite-temperature phase diagrams [56].
 However, this power comes at a substantial computational cost.
@@
 Determining a melting curve requires sampling thousands of large-scale atomic configurations through molecular dynamics, a task that has remained prohibitively expensive within DFT+DMFT.
-Recently, machine learning has begun to be explored as a way to reduce the cost of DMFT calculations, for example through surrogate impurity solvers, learned corrections to approximate Green’s functions, and data-driven models trained on simplified parameter spaces [mlgfMitra2025, mlgfDong2024, mlgfSturm2021, mlgfArsenault2014].
+Recently, machine learning has begun to be explored as a way to reduce the cost of DMFT calculations, for example through surrogate impurity solvers, learned corrections to approximate Green’s functions, and data-driven models trained on simplified parameter spaces [51, 20, 66, 6, 73, 53].
 While these studies demonstrate that key DMFT quantities can often be approximated efficiently in selected settings, most existing approaches have been confined to model Hamiltonians or narrow material classes, and the learned targets are not always constructed to respect the analytic structure and symmetry constraints required in realistic DFT+DMFT workflows.
 It therefore remains unclear whether machine learning can provide a robust, physics-consistent acceleration of DFT+DMFT for correlated solids with the accuracy needed for thermodynamic predictions, particularly under the extreme conditions relevant to Earth’s core.
-In this work, we demonstrate that embedding physical constraints into machine-learned self-energies enables, for the first time, systematic DFT+DMFT determination of the iron melting curve at core conditions.
-Our approach exploits the analytic structure of the self-energy: we decompose it into a static high-frequency limit $\Sigma(\infty)$ and a frequency dependent term $\Sigma(i\omega)$, represented compactly through Legendre coefficients, with each component predicted by E(3)-equivariant graph neural networks that respect the point-group symmetry of the local atomic environment.
-This physics-informed representation reduces the learning problem to smooth, real-valued targets while preserving the causal structure essential for DMFT self-consistency [mlgfArsenault2014].
+In this work, we demonstrate that embedding physical constraints into machine-learned self-energies enables scalable finite-temperature DFT+DMFT thermodynamics for correlated materials, and we use this capability to determine the hcp-Fe melting curve at core conditions as a demanding application.
+Our approach exploits the analytic structure of the self-energy: we decompose it into a static high-frequency limit $\Sigma(\infty)$ and a frequency-dependent term $\Sigma(i\omega)$, represented compactly through Legendre coefficients, with each component predicted by E(3)-equivariant graph neural networks that respect the point-group symmetry of the local atomic environment.
+This physics-informed representation reduces the learning problem to smooth, real-valued targets that respect the analytic structure required by the DFT+DMFT self-consistency cycle [6].
 The resulting model accelerates DMFT convergence by a factor of two to four across Fe, FeO, and NiO, enabling high-throughput generation of correlated electronic structure data.
-We leverage this capability to train a machine learned interatomic potential (MLIP) for iron and perform two-phase coexistence simulations spanning 307–352 GPa.
-Our predicted melting temperature of 6225K at 330 GPa agrees well with the recent experimental measurements [emcAnzellini2013, emcKraus2022, emcBalugani2024], providing strong validation that large-scale molecular dynamics (MD) simulations using DMFT-based MLIPs are viable for phase-equilibrium calculations in planetary interiors.
+We leverage this capability to train a machine learned interatomic potential (MLIP) for iron and perform solid–liquid coexistence simulations in the NVE ensemble [52, 3] spanning 307–352 GPa.
+Our predicted melting temperature of 6225K at 330 GPa agrees well with the recent experimental measurements [5, 40, 7], providing strong validation that large-scale molecular dynamics (MD) simulations using DMFT-based MLIPs are viable for phase-equilibrium calculations in planetary interiors.
 Within DFT+DMFT, electronic correlations enter entirely through the local self-energy.
@@
 First, the DMFT self-energy is complex-valued and, when obtained from quantum Monte Carlo impurity solvers, can be noisy at high frequencies.
-Second, $\Sigma(i\omega_{n})$ obeys strict analytic constraints: it is a causal function with a well-defined high-frequency expansion, $\Sigma(i\omega_{n})=\Sigma(\infty)+{\cal O}(1/i\omega_{n})$ [sigmaWang2011], and its imaginary part has definite sign on the imaginary axis.
+Second, $\Sigma(i\omega_{n})$ obeys strict analytic constraints: it is a causal function with a well-defined high-frequency expansion, $\Sigma(i\omega_{n})=\Sigma(\infty)+{\cal O}(1/i\omega_{n})$ [76], and its imaginary part has definite sign on the imaginary axis.
 Direct regression on a large set of complex values risks violating these constraints unless they are enforced explicitly.
@@
 The real part of the self-energy approaches a frequency-independent limit as $\omega_{n}\rightarrow\infty$:
-$Re\left\{\Sigma(i\omega_{n})\right\}\xrightarrow{\omega_{n}\to\infty}\Sigma(\infty),$ (1)
+[…]
 while the imaginary part decays to zero. $\Sigma(\infty)$ captures the static Hartree-Fock contribution from local interactions. We therefore decompose the self-energy as
@@
 This separation isolates the dynamical correlations in $\Delta\Sigma$ from the static component $\Sigma(\infty)$, each of which can be learned independently.
-The dynamical part $\Delta\Sigma(i\omega)$ is then Fourier transformed to imaginary time and expanded in a Legendre polynomial basis [Boehnke2011],
+The dynamical part $\Delta\Sigma(i\omega)$ is then Fourier transformed to imaginary time and expanded in a Legendre polynomial basis [13],
 $\Delta\Sigma(\tau)=\sum_{\ell=0}^{\ell_{\max}}\sqrt{2\ell+1}\,P_{\ell}(x(\tau))\,\Sigma_{\ell},$ (3)
@@
 where $\Delta\Sigma^{\mathrm{ML}}(i\omega_{n})$ is obtained by transforming the Legendre expansion back to the Matsubara axis.
-Both forward and reverse transformations are performed by the TRIQS software package [Parcollet2015, Boehnke2011].
+Both forward and reverse transformations are performed by the TRIQS software package [55, 13].
 This $\Sigma^{\mathrm{ML}}$, together with the predicted $E_{f}$, provides the initial condition for the DMFT self-consistency loop.
-Including $E_{f}$ as a learned quantity is an essential for achieving substantial speedups.
+Including $E_{f}$ as a learned quantity is essential for achieving substantial speedups.
 In DFT+DMFT the Fermi level is normally adjusted iteratively so that the total electron number satisfies the charge constraint, and poor initial guesses can severely delay convergence.
 By treating $E_{f}$ on the same footing as $\Sigma(\infty)$ and the Legendre coefficients, we define a joint functional $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ that maps a crystal structure to a complete, physics-consistent set of input parameters.
-To implement this functional, we represent each crystal structure as a periodic atomic graph and approximate $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ with an $E(3)$-equivariant graph neural network built using the open-source e3nn software package [e3nnGieger2022, e3nnKondor2018, e3nnThomas2018, e3nnWeiler2018].
+To implement this functional, we represent each crystal structure as a periodic atomic graph and approximate $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ with an $E(3)$-equivariant graph neural network built using the open-source e3nn software package [25, 37, 71, 77].
 Atoms are nodes, connected to neighbors within a radial cutoff, and are encoded by learned embeddings of their chemical species and mass.
@@
 To assess the accuracy of the physics-informed functional, we first compare its predictions for the self-energy against fully converged DFT+DMFT results on held-out test structures.
-For Fe and FeO, structural configurations were generated from MD simulations of high‑symmetry crystalline prototypes (e.g., hcp/bcc/fcc for Fe, B2 phase for FeO) to ensure coverage of both ordered and disordered environments.
-2.5‑ps DFT‑MD trajectories (0.5 fs timestep) of 4–8 atom cells were generated and representative snapshots were selected by k‑medoids clustering in structural fingerprint space [fpZhu2016], yielding 500 configurations in total (100-200 per trajectory).
+For Fe and FeO, structural configurations were generated from MD simulations of high-symmetry crystalline prototypes (e.g., hcp/bcc/fcc for Fe, B2 phase for FeO) to ensure coverage of both ordered and disordered environments.
+2.5-ps DFT-MD trajectories (0.5 fs timestep) of 4–8 atom cells were generated and representative snapshots were selected by k-medoids clustering in structural fingerprint space [80], yielding approximately 500 configurations per compound (100–200 per seed trajectory).
 Since the fingerprint represents the chemical environment of each crystal structure, the fingerprint distance allows us to quantify the overall difference in chemical environments between any two crystal structures.
 This ensures that we obtain diverse chemical environments for the solver to “learn” in an attempt to aid generalization of results to unseen structures.
-For NiO we use configurations from our previous study [Rao2024].
+For NiO we use configurations from our previous study [60].
 Full details of dataset generation and selection are provided in the Supplemental Material.
@@
 2 iterations).
-Since each CTQMC iteration dominates the computational cost, this translates directly into a factor of 2–4 reduction in wall-clock time per structure.
+Since CTQMC iterations dominate the computational cost, this translates directly into a factor of 2–4 reduction in wall-clock time per structure.
 The learned functional thus acts as a nearly cost-free preconditioner for the DMFT loop, turning the self-energy from a repeated, expensive output into a one-shot, physics-informed initial condition.
+FeO retains the largest iteration count under ML initialization, reflecting the complex Fe 3d3d–O $2p$ hybridization in the B2 phase and the correspondingly stronger frequency dependence of the self-energy, which makes the DMFT loop more sensitive to residual errors in the higher-order Legendre coefficients (see Supplemental Material).
 Leveraging this acceleration, we performed high-throughput DFT+DMFT calculations on an additional 600 Fe configurations at pressures and temperatures relevant to Earth’s core, increasing the Fe dataset to 1100 configurations with correlated energies and forces.
-Using all Fe data collected so far, we trained a NequIP-based machine-learned interatomic potential (MLIP) [nequipTan2025, nequipBatzner2022], achieving test-set mean absolute errors of 69.2 meV/atom for the energy and 76.7 meV/Å for the forces.
-Because the reference energies and forces are obtained from a CTQMC impurity solver, the training labels carry finite statistical uncertainty [huangMachineLearningDiffusion2023]; accordingly, the reported test errors reflect both model discrepancy and an intrinsic noise floor.
+Using all Fe data collected so far, we trained a NequIP-based machine-learned interatomic potential (MLIP) [69, 8], achieving test-set mean absolute errors of 69.2 meV/atom for the energy and 76.7 meV/Å for the forces.
+Because the reference energies and forces are obtained from a CTQMC impurity solver, the training labels carry finite statistical uncertainty [34]; accordingly, the reported test errors reflect both model discrepancy and an intrinsic noise floor.
 In the high-temperature regime of interest (e.g., $T=5000$ K, $k_{B}T\approx 430$ meV), the resulting per-atom energy error remains sub-$k_{B}T$, enabling stable large-scale MD and solid–liquid coexistence simulations.
-To assess sensitivity to the electronic temperature used for DFT+DMFT labeling, we recomputed the electronic free energies for representative Fe configurations at T = 5000 and 6000 K and find changes in relative free-energy differences of $\leq$ 6 meV/atom, well below the MLIP test errors.
-We determined melting temperatures using the solid–liquid coexistence method [aimcAlfe2009], which avoids the hysteresis inherent in single-phase superheating or supercooling approaches.
+The per-atom global energy MAE does not directly determine the melting-temperature error; the relevant quantity is the phase-distinguishing signed residual in the solid–liquid energy/free-energy difference.
+A phase-resolved analysis on the held-out Fe test set gives a central $\delta T_{m}$ estimate of $\sim 20$ K, while alternative phase partitions give a conservative systematic envelope of order $10^{2}$ K, reported separately from the $\pm 42$ K internal coexistence fluctuation (see Supplemental Material for the full propagation, including the $L_{\mathrm{melt}}=T_{m}\,\Delta S_{m}$ derivation [2, 3]).
+The MLIP force MAE of 76.7 meV/Å is comparable to or lower than values reported in recent finite-temperature and extreme-condition MLIP studies, including melting and crystallization applications [18, 17, 22, 59, 70, 78], supporting the fidelity of the coexistence trajectories.
+Residual stress anisotropy of 2–3 GPa from the fixed-cell NVE protocol contributes a separate pressure-axis uncertainty (see Supplemental Material).
+To assess sensitivity to the labelling temperature we recomputed representative relative free-energy differences at 5000 and 6000 K and find changes of $\leq 6$ meV/atom, well below the MLIP test errors.
+We determined melting temperatures using solid–liquid coexistence simulations in the NVE ensemble [52, 3], which avoid the hysteresis inherent in single-phase superheating or supercooling approaches.
 An elongated supercell (9216 atoms) containing coexisting solid (hcp) and liquid regions is equilibrated within the NVE ensemble; if the temperature lies above $T_{m}$, the solid region melts, whereas below $T_{m}$ the liquid crystallizes.
-The phase fractions were monitored using local bond-order analysis [pfaLechner2008].
+The phase fractions were monitored using local bond-order analysis [46].
 Figure 2 illustrates a representative coexistence configuration at 307.8 GPa: the density profile along the long axis exhibits pronounced oscillations in the crystalline region, characteristic of atomic layering, and a comparatively smooth profile in the liquid region, confirming a well-defined solid–liquid interface.
@@
 Pressure was sampled from finite-difference stress calculations including the kinetic contribution, and averaged over the final 5 ps.
+To verify that the MLIP remains well-behaved in the disordered interface region, we rely on two diagnostics: the training set already contains liquid-like and high-temperature solid configurations sampled from DFT-MD seeds run at or above the small-cell melting temperature, and the coexistence MD itself shows stable behaviour—over tens of picoseconds the interface remains stationary, the NVE energy drift is consistent with integration error rather than with model instability, and the interfacial density profile is smooth (see Supplemental Material).
 Figure 2: Density profile along the long axis of simulation cell taken at 20 ps into equilibration for simulation at 307.8 GPa.
 Shown in the insets are a snapshot of the atomic configuration at 20 ps as well as the temperature variation over the length of the simulation, which stabilizes around 15 ps.
-The 2 coexistence interfaces can been seen on both sides of the liquid.
-Our melting results, spanning 307–352 GPa, is shown in Fig.
-3 alongside experimental determinations and prior ab initio predictions.
-The results are well described by a Simon-Glatzel relation [simonglatz],
+The 2 coexistence interfaces can be seen on both sides of the liquid.
+Our melting results, spanning 307–352 GPa, are shown in Fig. 3 alongside experimental determinations and prior ab initio predictions. The results are well described by a Simon-Glatzel relation [11],
 $T_{m}=T_{0}\left(\frac{P-P_{0}}{a}+1\right)^{1/c},$ (5)
 with $T_{0}=6167$ K, $P_{0}=323.1$ GPa, $a=298.73$ GPa, and $c=2.43$.
-At the inner-core boundary pressure of 330 GPa, we obtain $T_{m}=6225\pm 42$ K, in excellent agreement with the laser-heated diamond anvil cell measurements of Anzellini et al.
-($6230\pm 500$ K) [emcAnzellini2013] and the ramp-compression experiments of Kraus et al.
-[emcKraus2022].
-It is also compatible with the recent shock-compression XAS constraints of Balugani et al., who extrapolate an upper bound of 6202 ± 514 K for Fe at 330 GPa [emcBalugani2024].
-Our result lies above the resistance-heated DAC determination of Sinmyo et al.
-($5500\pm 220$ K) [emcSinmyo2019]; however, as noted in Ref.
-[emcAnzellini2013], different melting criteria in static experiments can yield systematically different temperatures, and the discrepancy among experiments themselves remains an active area of investigation.
+At the inner-core boundary pressure of 330 GPa, we obtain $T_{m}=6225\pm 42\,\mathrm{K}_{\mathrm{stat}}\pm\mathcal{O}(10^{2})\,\mathrm{K}_{\mathrm{sys}}$, where the first bar reflects the internal fluctuation of the coexistence MD over the final 5 ps and the second is the systematic uncertainty from the phase-distinguishing component of the MLIP error described above (Supplemental Material).
+This is in agreement with the laser-heated diamond anvil cell measurements of Anzellini et al.
+($6230\pm 500$ K) [5] and the ramp-compression experiments of Kraus et al.
+[40], both of which rely on extrapolation to the ICB (upward from $\sim$200 GPa and downward from $\sim$600 GPa, respectively), so that the quoted experimental bars themselves already include substantial extrapolation uncertainty.
+Our result is also compatible with the recent shock-compression XAS constraints of Balugani et al., who extrapolate an upper bound of 6202 $\pm$ 514 K for Fe at 330 GPa [7].
+It lies above the resistance-heated DAC determination of Sinmyo et al.
+($5500\pm 220$ K) [62]; however, as noted in Ref.
+[5], different melting criteria in static experiments can yield systematically different temperatures, and the discrepancy among experiments themselves remains an active area of investigation.
 Our agreement with recent experimental constraints also provides insight into why melting predictions for Fe at core pressures have remained so dispersed.
-Most published ab initio melting curves based on standard DFT span a broad range and are frequently higher than experimental estimates, with discrepancies reaching several hundred kelvin and an overall scatter approaching $10^{3}$ K [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009].
-This sensitivity likely reflects, in part, that commonly used high-$T$ DFT melting workflows dynamical correlations only approximately, thereby missing high-temperature spin fluctuations that contribute to the free energy [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015].
-Fe at these conditions is paramagnetic, yet thermal spin fluctuations can sustain sizable instantaneous local moments and associated entropy contributions [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015].
+Most published ab initio melting curves based on standard DFT span a broad range and are frequently higher than experimental estimates, with discrepancies reaching several hundred kelvin and an overall scatter approaching $10^{3}$ K [3, 26, 68, 67, 10, 65, 2, 9, 79, 63].
+This sensitivity likely reflects, in part, that commonly used high-$T$ DFT melting workflows incorporate dynamical correlations only approximately, thereby missing high-temperature spin fluctuations that contribute to the free energy [61, 57, 74].
+Fe at these conditions is paramagnetic, yet thermal spin fluctuations can sustain sizable instantaneous local moments and associated entropy contributions [61, 57, 74].
 If these magnetic/electronic excitations are more pronounced in the liquid than in the solid, they increase the entropy of melting and reduce the coexistence temperature through the equilibrium condition $\Delta G(P,T_{m})=0$, i.e.
@@
 DFT+DMFT addresses this limitation by treating the $3d$ correlations dynamically through the frequency-dependent self-energy $\Sigma(i\omega_{n})$, capturing quasiparticle renormalization and fluctuating local moments on equal footing.
-That our melting curve falls within experimental uncertainty while using interaction parameters $(U,J)$ adopted from prior DFT+DMFT studies [ieccPourovskii2017, chsGavriliuk2023, chsHo2024] (rather than fitted to melting data) indicates genuine predictive power and is consistent with the view that dynamical correlations account for a substantial part of the long-standing discrepancy between simulation and experiment for Fe at core conditions.
-Figure 3: Full melting curve of Iron as predicted by our 2-phase simulations and compared to data from various experimental [emcAnzellini2013, emcSinmyo2019, emcLi2020, emcKraus2022, emcBalugani2024] computational studies [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009].
-In conclusion, we introduced a physics‑informed machine‑learning approach that substantially reduces the computational cost of realistic DFT+DMFT calculations.
-Rather than directly fitting the noisy Matsubara self‑energy as an unconstrained output, we learn a compact representation built from the high‑frequency limit and a truncated set of real Legendre coefficients, together with the Fermi level.
-Implemented with E(3)‑equivariant graph neural networks, this provides a physically consistent warm start to the DMFT fixed point and reduces the number of costly impurity iterations by a factor of two to four across Fe, FeO, and NiO.
+That our melting curve falls within experimental uncertainty while using interaction parameters $(U,J)$ adopted from prior DFT+DMFT studies [57, 23, 33] (rather than fitted to melting data) indicates genuine predictive power and is consistent with the view that dynamical correlations account for a substantial part of the long-standing discrepancy between simulation and experiment for Fe at core conditions.
+Figure 3: Full melting curve of Iron as predicted by our solid–liquid coexistence simulations and compared to data from various experimental [5, 62, 48, 40, 7] computational studies [3, 26, 68, 67, 10, 65, 2, 9, 79, 63].
+In conclusion, we introduced a physics-informed machine-learning approach that substantially reduces the computational cost of realistic DFT+DMFT calculations.
+Rather than directly fitting the noisy Matsubara self-energy as an unconstrained output, we learn a compact representation built from the high-frequency limit and a truncated set of real Legendre coefficients, together with the Fermi level.
+Implemented with E(3)-equivariant graph neural networks, this provides a physically consistent warm start to the DMFT fixed point and reduces the number of costly impurity iterations by a factor of two to four across Fe, FeO, and NiO.
 This acceleration turns DFT+DMFT into a practical engine for generating correlated energies and forces at scale.
-Using the accelerated workflow, we expanded our Fe dataset and trained a NequIP‑based MLIP that enables large‑scale molecular dynamics for phase equilibrium.
-Applying solid–liquid coexistence simulations in the 307–352 GPa range, we obtained the melting curve of Fe under core conditions and predict $T_{m}=6225$ K at 330 GPa, in close agreement with recent experimental constraints.
-More broadly, this approach turns DFT+DMFT into a scalable engine for finite‑temperature simulations of correlated materials, with applications ranging from transition‑metal compounds to matter under planetary interior conditions.
+Using the accelerated workflow, we expanded our Fe dataset and trained a NequIP-based MLIP that enables large-scale molecular dynamics for phase equilibrium.
+Applying solid–liquid coexistence simulations in the 307–352 GPa range, we obtained the hcp-Fe melting curve under core conditions and predict $T_{m}=6225$ K at 330 GPa, in close agreement with recent experimental constraints.
+More broadly, the present framework provides a route toward scalable finite-temperature simulations of correlated materials, with applications ranging from transition-metal compounds to matter under planetary interior conditions.
+In this work, the learned self-energy is used strictly as a warm start, and all correlated reference energies and forces are obtained after full DFT+DMFT refinement.
+Nevertheless, the near-converged character of the ML initialization suggests a future uncertainty-controlled surrogate mode, in which full CTQMC refinement is invoked selectively rather than for every configuration, potentially enabling substantially larger speedups.
 Acknowledgements—This work was supported by the startup funds of the office of the Dean of SASN of Rutgers University-Newark.
@@
 Data Availability—The software used in prediction of the self-energies and Fermi level has been made available at https://github.com/Rutgers-ZRG/SigML
+Supplemental Material
+S1 Neural Network Structure and Training
+Our approach utilizes neural networks that are equivariant to rotations and inversion, and invariant to translations.
+The model is implemented using the open-source e3nn framework [24], and operates on periodic atomic graphs constructed from crystalline input.
+Each atom is represented as a node in the graph, and edges are formed between neighboring atoms within a fixed radial cutoff.
+These edges encode the relative position vectors, enabling the use of spherical harmonics $Y_{l}(r)$ and learned radial basis functions to construct equivariant edge features.
+Each atom is encoded via mass-weighted and atomic number one-hot vectors, which are passed through a shared embedding layer to produce a scalar feature representations.
+These embedded features are used as the input node features and node attributes.
+The resulting periodic graph is processed by a sequence of convolution and gated nonlinearity layers.
+Each layer consists of an equivariant convolutional operation followed by a gated non-linear activation function, where scalar channels modulate higher-order spherical tensor channels.
+We use 2 interaction layers with an embedding dimension of 64 and a maximum irrep order of 2.
+This structure preserves full equivariance throughout the network while allowing expressive nonlinearity.
+[…]
+After several layers of message passing, a final convolution is applied without gating.
+The resulting $l=0$ features are passed through a parametric ReLU (PReLU) nonlinearity.
+To reduce complexity of the output, we predict the self-energy for only the symmetrically inequivalent atoms.
+To predict the self-energy, we adopt an ensemble approach.
+Specifically, we instantiate an independent network for each orbital, sharing the same structure but maintaining distinct learnable weights.
+The outputs from each orbital-specific network are stacked to form the final output, which is set to the Legendre coefficients for the specific orbital self-energy.
+For the fermi-level prediction, we simply sum the resulting $l=0$ features with a single instantiation of the network.
+The resulting models were trained for 20 epochs using an AdamW optimizer [49] and a learning rate of 0.01, adaptively tuned using a learning rate scheduler based on validation loss.
+The loss was calculated as the smooth L1 loss as implemented in PyTorch, with a beta value of 1.
+S2 DFT+DMFT Calculation Details
+The DFT part of the DFT+DMFT loop was carried out using the Local Density Approximation (LDA) with the augmented plane wave plus local orbital method, as implemented in the WIEN2K package [12].
+For Fe, a muffin-tin radius of 1.6 bohr was used along with a k-point mesh proportional to the number of Fe atoms in the unit cell.
+Here we use 2000/(# of Fe atoms).
+The RKMax value was fixed at 7.5 and an energy cutoff of -7.5 Rydberg was chosen to separate core and valence states.
+DFT calculations were iterated until charge convergence of 0.00001 was achieved.
+For FeO, a muffin-tin radius of 1.6 bohr for Fe and 1.4 bohr for O was used along with a k-point mesh proportional to the number of atoms in the unit cell.
+The same scheme for choosing the number of k-points was used as the Fe calculations.
+The RKMax value was fixed at 7.5 and an energy cutoff of -7.5 Rydberg was chosen to separate core and valence states.
+DFT calculations were iterated until charge convergence of 0.00001 was achieved.
+For NiO, a muffin-tin radius of 1.8 bohr for Ni and 1.5 bohr for O was used.
+An LDA exchange correlation functional was used with a k-point mesh of 1000 k-points.
+The RKMax value was fixed at 7.5 and an energy cutoff of -6 Rydberg was chosen to separate core and valence states.
+DFT calculations were iterated until charge convergence of 0.001 was achieved.
+The DMFT part of the DFT+DMFT loop was handled by the eDMFT software package [30, 28].
+The eDMFT package utilizes a hybridization expansion of the impurity action, which is subsequently sampled using monte carlo.
+Calculations are carried out at a temperature of 5000 K for Fe and FeO, and at 611 K for NiO.
+Coulomb interaction values of 5.0 eV (Fe), 10.0 eV (FeO), or 8.0 eV (NiO) in the density-density form and Hund’s coupling values of 0.93 eV (Fe), 1.0 eV (FeO), and 0.9 eV (NiO) were used to build the impurity Hamiltonian.
+These values were chosen based on previous studies [57, 23, 33] and are constant across all calculations.
+The energy window for construction of the DMFT projector was chosen to span from -10 to 10 eV for NiO and FeO, and -12.2 to 4 eV for Iron, around the Fermi level.
+The impurity problem was then solved using the continuous time quantum monte carlo (CTQMC) solver as implemented by eDMFT [31] with 108 MC steps per processor for 32 processors for Fe and FeO, and 107 MC steps per processor for 32 processors for NiO.
+Calculations were stopped when $|n_{imp}-n_{lat}|\leq 0.01$, where $n_{imp}$ and $n_{lat}$ are the impurity and lattice occupation numbers respectively.
+Forces are also calculated for each structure using the derivative of the Luttinger-Ward functional [29].
+Double counting is handled using a nominal double counting scheme, as introduced in Refs [30, 27].
+We treat the transition metal atoms as correlated, using a stationary projector based on the solution of the Schrodinger equation within the LDA muffin-tin sphere.
+Spin-orbit coupling is ignored for all compounds because spin-orbit coupling is expected to be weak for 3d transition metal atoms.
+We treat all systems in the paramagnetic state (no long-range magnetic order / no spin-polarized DFT); local moment fluctuations are included at the DMFT level.
+S3 Training set generation details
+For Fe and FeO, we perform MD simulations using the Vienna ab-initio simulation package (VASP) [43, 41, 42, 44] at high temperatures in order to sample a large area of configuration space.
+We sample configurations from both before and after melting in order to capture the solid and liquid phase behaviors, as the goal is a MLIP that can produce accurate predictions for both phases.
+For Fe, we perform MD simulations with the HCP ($\epsilon$), BCC ($\alpha$), and FCC ($\gamma$) structures as a starting point.
+For each phase, we set pressure to 300 GPa and allow cell parameters and volume to vary for 2.5 ps with a 0.5 fs timestep at a start temperature of 6000 K.
+From these configurations, we use the fingerprint clustering described above to select out 100-200 structures each.
+Our final dataset consists of 500 structures, as unconverged calculations were not included in the training set.
+In order to enhance the MLIP, 2 more datasets were generated for Fe.
+One dataset was generated from MD simulations of the $\epsilon$-Fe phase taken at 20 different volumes ranging from volumes of 8.3 Å/atom to 6.5 Å/atom, within the NVT ensemble in order to capture high pressure behavior of Iron.
+The MD simulations were run at a temperature of 5, [^0] K.
+These calculations were then clustered according to their relative fingerprint distances, with 10 points being selected for each volume.
+The next dataset was generated using the M3GNET [15] universal interatomic potential by running MD simulations at 10, [^0] K within the NPT ensemble, with pressure being targeted at 330 GPa in order to capture the melting behavior near the inner-core boundary (ICB).
+200 structures were selected from the solid and liquid phases each using fingerprint clustering once again.
+The self-energy and Fermi level were then predicted for each new structure, and used as a starting condition for the DFT+DMFT loop.
+These 600 structures with their respective energies and forces were added to the total Fe DFT+DMFT-labeled configuration pool, bringing the total to 1100 configurations; after removal of nonconverged calculations 1052 unique structures were used for MLIP training/validation/testing (see below).
+For FeO, MD simulations were carried out for the B2 ($pm\overline{3}m$) phase at temperatures of 10, [^0] K, as this is believed to be the dominant phase at conditions near the ICB [54].
+Simulations were run for 2.5 ps with a 0.5 fs timestep within the NVT ensemble for 1 unit cell of FeO, from which 500 structures were selected through fingerprint clustering.
+For NiO, we utilize training data from a previous study on the melting curve [60].
+Specifically, The training data was generated by performing 500 random perturbations of the atomic positions and lattice constants for the trigonal, face-centered cubic and body centered cubic phases each for a total of 1500 training structures.
+For each structure, the volume was also randomly perturbed, down to 65% of the equilibrium volume, in order to capture behavior at high pressures.
+After the volume change, perturbations to the structure include displacement of the unit cell basis vector lengths by a random amount between 0.3 Å and -0.3 Å, then a change in the angles by a random amount between 22.5 to -22.5 degrees, and displacement of the atoms in the unit cell from their original positions by a random value between -10% to 10% of each of the newly changed unit cell vectors.
+The clustering in fingerprint space of the Fe and FeO MD trajectories serves to create a diverse but manageable, training set.
+Since the fingerprint represents the chemical environment of each crystal structure, the fingerprint distance allows us to quantify the overall difference in chemical environments between any two crystal structures.
+This ensures that we obtain diverse chemical environments for the solver to "learn" in an attempt to aid generalization of results to unseen structures.
+Shown in Fig S1 are the results of the clustering within a 3D space generated by first 3 components of a principal component analysis decomposition of the structural fingerprints for Fe.
+The arcs within fingerprint space show the changing chemical environment generated by the molecular dynamics trajectories.
+Crucially, the k-medoids algorithm is able to cluster effectively using the fingerprint distance matrix.
+Figure S1: Trajectories within PCA decomposed fingerprint space for Fe structures generated using molecular dynamics.
+The 3 axes represent the first 3 components of the PCA.
+The colors represent structures considered to be within the same cluster.
+Cluster centers are shown by larger dots.
+S4 Training of the Machine Learned Interatomic Potential
+We utilize the open-source NequIP package [69, 8] in order to construct E3-equivariant machine learned interatomic potentials.
+We train on 1,052 unique Fe structures, with 48 structures removed from the total set due to nonconverged calculations.
+We reserve 10% of the dataset for testing, 10% for validation and train on the remaining 80%.
+Our interatomic potential is constructed using a cutoff radius of 5 Å and contains 4 interaction blocks with 32 features.
+We use a maximum irrep order of 2, with 2 radial layers each with 64 neurons.
+For the basis, we use 8 radial Bessel functions with a polynomial cutoff of 6.
+We also utilize a ZBL pair potential [81] for extremely close range interactions in order to reduce erroneous results during MD.
+Shown in Figure S2(a) and (b) are the errors for the energy and force model respectively on the test set of Fe structures.
+We achieve an overall mean absolute error for energy of 69.2 meV/atom and an overall mean absolute error for force of 76.7 meV/Å.
+Figure S2: Predicted versus actual (a) energies and (b) force components. Results are displayed for the 105 structures in the Fe testing set, with the black dashed line representing perfect agreement.
+S5 Validation of pressure-volume curve
+Presented in Figure S3 is our predicted pressure versus volume curve as compared to experimental data.
+The Vinet equation [75] was used along with corrected fit parameters provided in Ref [32].
+We find good agreement in the pressure range around the ICB pressure of around 330 GPa.
+Since our training set is specialized around this pressure, we see divergences from the experimental results at much higher and lower pressures.
+Figure S3: Pressure versus volume for our MLIP compared to Vinet equation fit of experimental data from Refs [32, 50, 21, 19].
+Experimental pressure measurement data was corrected according to the treatment in Ref [32] using an updated Pt scale.
+S6 Melting Point Simulation Details
+Melting point simulations were carried out using solid–liquid coexistence in the NVE ensemble [52, 3] using the atomic simulation environment software package [45].
+Follow standard practices, we first thermalize the system near the melting point within the NVT ensemble using Langevin dynamics.
+From there, we freeze the solid half the simulation cell and raise the temperature, allowing the liquid half to melt.
+Once full melt is established, the liquid phase is brought back down to near the predicted melting temperature, and the entire system is rethermalized.
+At that point, the ensemble is switched to NVE and solid liquid phases are allowed to interact.
+The phase boundary is monitored using local bond-order analysis [46] and melting is also confirmed visually by examining the simulation cells.
+Simulations are performed with 9216 Fe atoms within a nearly orthorhombic unit cell created from the hcp-Fe structure.
+The lattice parameters for the hcp-Fe structure were obtained from the experimental data of Ref [32] and optimized using finite differences stress for various pressures.
+Simulations are allowed to evolve for 20-30ps depending on phase stabilization and temperature and pressure data is collected over the last 5ps of simulation.
+Presented in table S1 is our raw data for predicted melting points versus sampled pressure within the simulation cell.
+Pressure was calculated using finite differences under deformations of the unit cell, with the contribution from temperature being included through the kinetic part of the virial stress tensor.
+In Figure S4, we show the phase fractions over the length of the simulation for each measured pressure point.
+Simulations stabilized around 15ps.
+We also attempt to quantify uncertainty by calculation of the standard deviation of pressure and temperature over the last 5ps of simulation.
+We present the results in Table S1.
+Errors in pressure are on the order of 1GPa, while errors in temperature are on the order of 50K.
+$P$ (GPa) $T_{m}$ (K) $\Delta P$ $\Delta T$
+307.8 6026.13 0.62 30.7
+323.1 6167.13 0.85 42.3
+337.2 6252.97 0.80 41.5
+352.4 6420.19 0.99 40.9
+Table S1: Pressure versus predicted melting temperature for $\epsilon$-Fe.
+$\Delta P$ and $\Delta T$ are the standard deviations of the instantaneous pressure and temperature over the final 5 ps of the NVE run, i.e.
+temporal fluctuations only; a separate, larger systematic pressure-anisotropy uncertainty of 2–3 GPa is discussed in Sec.
+S11.
+Figure S4: Phase fraction over the length of simulation for each measured pressure value. Most simulations stabilize around 15ps.
+S7 Legendre-coefficient decay across compounds
+The dynamical part of the self-energy $\Delta\Sigma(\tau)$ is expanded in Legendre polynomials with coefficients $\Sigma_{\ell}$ that decay with $\ell$ at a rate set by the smoothness of $\Delta\Sigma$ on the imaginary-time axis, and therefore by the frequency structure of the self-energy.
+For the three compounds studied, we find (i) for Fe at 5000 K, the coefficients fall below $10^{-2}$ in Hartree units by $\ell\approx 20$; (ii) for FeO at 5000 K in the B2 phase, the decay is markedly slower, with significant weight retained up to $\ell\approx 30$, consistent with the stronger Fe 3d3d–O $2p$ hybridization producing a more frequency-dependent self-energy; (iii) for NiO at 611 K, a finer Matsubara grid resolves lower-frequency structure that requires $\ell_{\max}\approx 70$ to reproduce accurately.
+This compound ordering of required $\ell_{\max}$ is consistent with the observed iteration counts in the main text: FeO retains the largest DMFT iteration count under ML initialization because a given tolerance on $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|$ is more sensitive to residual errors in a self-energy with stronger frequency dependence.
+S8 Propagation of MLIP error into the melting temperature
+The per-atom energy MAE of the Fe MLIP, $E_{\mathrm{MAE}}=69.2$ meV/atom, is a global diagnostic of the fit quality but is not identical to the residual that enters the coexistence melting temperature.
+The melting temperature is set thermodynamically by the equality of solid and liquid chemical potentials, and the relevant MLIP-error metric is the phase-distinguishing component
+$\sigma^{\mathrm{phase}}_{\Delta E}\;\equiv\;\left|\,\langle\delta E\rangle_{\mathrm{liquid}}-\langle\delta E\rangle_{\mathrm{solid}}\,\right|,$ (S1)
+where $\delta E_{i}=E_{i}^{\mathrm{MLIP}}-E_{i}^{\mathrm{DFT+DMFT}}$ is the per-atom signed energy residual on the held-out test set.
+We use $\sigma^{\mathrm{phase}}_{\Delta E}$ as a proxy for the MLIP contribution to the solid–liquid enthalpy/free-energy bias, treating pressure-axis errors separately (Sec.
+S11).
+Two contributions to $E_{\mathrm{MAE}}$ do not appear as a coherent solid–liquid bias unless they are inherited in a phase-selective way: (i) the statistical component inherited from the CTQMC impurity solver, which is uncorrelated between training snapshots; and (ii) any systematic component common to both phases, which cancels in the chemical-potential difference.
+Both therefore contribute to the global MAE but should not contaminate the phase-distinguishing residual unless the model has learned them in a phase-selective way; the empirical check is the signed phase-resolved residual itself.
+To quantify $\sigma^{\mathrm{phase}}_{\Delta E}$ directly, we used the 106 Fe configurations available in the held-out pool (Fig.
+S2).
+We computed the per-configuration signed residuals $\delta E_{i}$ and partitioned configurations into solid-like and liquid-like environments.
+The cleanest, data-driven partition uses $k$-means clustering (k=2k{=}2) in the two-dimensional feature space spanned by the per-config force RMS $F_{\mathrm{rms}}$ and $\Delta E_{\mathrm{above\ floor}}(V)$, the per-config energy minus a cold lower-envelope at the same atomic volume (Fig.
+S5, left and middle panels).
+The clustering yields a solid-like cluster of 73 configurations (mean $F_{\mathrm{rms}}\approx 3.7$ eV/Å, $\Delta E_{\mathrm{above\ floor}}\approx 80$ meV/atom) and a liquid-like cluster of 33 configurations (mean $F_{\mathrm{rms}}\approx 6.6$ eV/Å, $\Delta E_{\mathrm{above\ floor}}\approx 820$ meV/atom).
+The phase-mean signed residuals are $\langle\delta E\rangle_{\mathrm{solid}}=+1.8$ meV/atom and $\langle\delta E\rangle_{\mathrm{liquid}}=-0.1$ meV/atom, giving $\sigma^{\mathrm{phase}}_{\Delta E}=1.9$ meV/atom—substantially smaller than the 69.2 meV/atom global MAE.
+The per-phase signed-residual histograms (Fig.
+S5, right panel) show that the two distributions overlap heavily and are individually centred near zero, with widths of order the global MAE: the spread within each phase reflects per-configuration model residuals together with finite CTQMC label uncertainty (both contribute to the global MAE but do not contribute coherently to the phase-mean signed residual unless learned in a phase-selective way), while the small offset between the phase means is what survives to bias $T_{m}$.
+As a robustness check we also explored ad-hoc partitions based on $F_{\mathrm{rms}}$ and $\Delta E_{\mathrm{above\ floor}}$ thresholds, which give $\sigma^{\mathrm{phase}}_{\Delta E}$ in the range 2–34 meV/atom (median 13 meV/atom); the variation reflects the inherent ambiguity of the solid/liquid distinction in 4-atom training cells at high temperature.
+Figure S5: Phase-resolved MLIP-residual analysis on the held-out Fe test set (106 4-atom configurations).
+Left: configuration-energy versus per-atom volume; the cold lower-envelope (black line) is the floor used to define $\Delta E_{\mathrm{above\ floor}}(V)$.
+Middle: two-dimensional feature space ($\Delta E_{\mathrm{above\ floor}}$, $F_{\mathrm{rms}}$); $k$-means clustering with k=2k{=}2 separates 73 solid-like (blue) from 33 liquid-like (red) configurations.
+Right: per-phase distributions of the signed residual $\delta E=E_{\mathrm{MLIP}}-E_{\mathrm{DFT+DMFT}}$ for the production NequIP model.
+The two distributions overlap heavily and are individually centred near zero with widths of order the global MAE; the small offset between the phase means yields $\sigma^{\mathrm{phase}}_{\Delta E}=1.9$ meV/atom, which corresponds to a Clausius–Clapeyron melting-temperature shift of $\sim 19$ K.
+Inserting the central $k$-means value into the Clausius–Clapeyron-type relation
+$\delta T_{m}\;\simeq\;T_{m}\,\frac{\sigma^{\mathrm{phase}}_{\Delta E}}{L_{\mathrm{melt}}},$ (S2)
+we adopt $L_{\mathrm{melt}}$ for Fe at ICB conditions inferred from $L_{\mathrm{melt}}=T_{m}\Delta S_{m}$, using the entropy of melting reported by Alfè et al.
+near ICB pressures, $\Delta S_{m}\approx 0.9$$1.1\,k_{B}$/atom [2], together with $T_{m}\approx 6300$64006400 K [3], giving $L_{\mathrm{melt}}\approx 0.5$$0.6$ eV/atom (with the $P\Delta V$ contribution already contained in $\Delta H_{m}=T_{m}\Delta S_{m}$).
+For the conservative estimate below we use a broader latent-heat bracket $L_{\mathrm{melt}}=0.4$$0.8$ eV/atom.
+Using the central 0.50.50.60.6 eV/atom range gives $\delta T_{m}\approx 20$2525 K for the $k$-means phase-resolved residual $\sigma^{\mathrm{phase}}_{\Delta E}$; the broader 0.40.40.80.8 eV/atom bracket gives $\delta T_{m}\approx 14$2929 K.
+Across alternative phase partitions, the corresponding systematic envelope remains of order $10^{2}$ K, reaching a few $\times 10^{2}$ K for the most conservative partition/latent-heat combination.
+Treating the full global energy MAE of 69.2 meV/atom as a perfectly phase-correlated solid–liquid bias would instead give $\delta T_{m}$ of order $5\times 10^{2}$$10^{3}$ K.
+The phase-resolved signed-error analysis suggests that this worst-case limit is overly conservative for the held-out data: phase-common contributions cancel in the solid–liquid difference, while uncorrelated label noise is not expected to contribute coherently to the phase-mean bias unless learned in a phase-selective way.
+In the main text we therefore report the internal coexistence fluctuation, $\pm 42$ K, separately from a conservative systematic uncertainty envelope of order $10^{2}$ K associated with the possible phase-distinguishing MLIP error.
+S9 MLIP behaviour in the solid–liquid interface
+We assess MLIP transferability into the disordered solid–liquid interface region through two complementary diagnostics.
+First, the Fe DFT+DMFT-labeled configuration pool (1100 configurations; 1052 used for training/validation/testing after removing nonconverged calculations) already includes configurations drawn from DFT-MD seed trajectories run near or above the relevant high-temperature melting regime, so liquid-like disordered local environments are represented alongside high-temperature solid configurations.
+Second, the coexistence MD trajectories provide stability diagnostics that would expose pathological model behaviour in the interface region: over tens of picoseconds the interface remains stationary, the interfacial density profile is smooth, and the NVE energy drift is consistent with integration error rather than with model instability.
+The absence of spurious re-melting, anomalous density jumps, or anomalous energy drift across all reported state points is consistent with adequate interpolation by the MLIP in the interface region.
+A configuration-by-configuration fingerprint-distance analysis of interface atoms against the training set is a quantitative complement to these diagnostics that we are pursuing as a follow-up.
+S10 Per-phase breakdown of the Fe training set
+The Fe DFT+DMFT-labeled configuration pool (1100 configurations) is populated by seed trajectories from high-symmetry crystalline prototypes at finite temperature together with a pressure/temperature active-learning enhancement, so that both solid- and liquid-like local environments for hcp, bcc, and fcc Fe are included.
+The breakdown across phase origins is: approximately 160–170 configurations each from hcp, bcc, and fcc seed trajectories at 300 GPa (500 total selected by $k$-medoids from a larger pool); 200 configurations from volume-scan NVT runs on hcp Fe at 5000 K spanning $V\in[6.5,\,8.3]$ Å3/atom; and 400 configurations selected from M3GNet-driven NPT MD at 10000 K near 330 GPa to include both solid and liquid environments near the ICB.
+In particular, bcc configurations are included in the training set.
+S11 Stress anisotropy in the NVE coexistence calculations
+Because the NVE coexistence ensemble fixes the supercell shape during MD, the three diagonal components of the stress tensor are not constrained to be equal, and some anisotropy of the stabilized stress is possible.
+For each reported state point we evaluated, over the final 5 ps of the NVE run, the deviation of each diagonal stress component from the mean of the three; the largest such deviation defines a per-state-point anisotropy.
+Averaged over our four state points this anisotropy is approximately 1.3 GPa, with a spread of 2–3 GPa across the diagonal components at the most anisotropic state points.
+The off-diagonal components of the stress tensor are negligible (below 0.01 GPa for the orthorhombic cells used here).
+This pressure-axis uncertainty is separate from—and larger than—the temporal trace-pressure fluctuations $\Delta P\sim 0.6$11 GPa tabulated in Table S1: $\Delta P$ is the standard deviation of the instantaneous trace-pressure over time at fixed cell shape, while the 2–3 GPa anisotropy quantifies the non-hydrostaticity of the stabilized stress tensor itself.
+The two contributions are reported separately when interpreting the pressure coordinate of the melting curve.